# Zhou, Orfanoudaki, and Zhu

### v2

@@ 4.2.3 Proposed Algorithm
where the radius $R(\alpha)$ is obtained by calibrating the non-conformity scores $\{L_{i}\}_{i=1}^{n}$ on $\mathcal{D}$. Of note, smaller $\alpha$ corresponds to larger $R(\alpha)$ and higher empirical coverage. We consider three specifications of the radius function $R(\alpha):[0,1]\to\mathbb{R}_{+}\cup\{\infty\}$, each satisfying the boundary conditions $R(\alpha)=\infty$ for $\alpha\in[0,1/(n+1))$ and $R(1)=0$. For $\alpha\in[1/(n+1),1)$, we specify:
@@
$R_{\infty}(\alpha)$ $=\infty,$ (Monte Carlo radius) (10)
for $\alpha\in[1/(n+1),1)$, where $\mathbbm{1}\{\cdot\}$ is the indicator function. Here, $R_{p}$ corresponds to the classical conformal radius based on the empirical quantiles of calibration residuals (vovk2005algorithmic, singh2024distribution); $R_{e}$ is an $e$-value–based variant (grunwald2024safe), offering stronger post-hoc validity guarantees (vovk2025conformal, balinsky2024enhancing, gauthier2025values); and $R_{\infty}$ is used in a Monte Carlo-based estimation. We will later discuss their respective validity properties, showing that $R_{e}$ provides the strongest robustness, $R_{\infty}$ the highest accuracy, and $R_{p}$ a balanced trade-off between the two.
@@
Comparative illustration of risk estimation using a point model and generative models ($K=3$ and $K=4$) in CREDO. The gray shaded area represents the inverse feasible region $\pi^{-1}(z)$. The blue dotted ellipsoids represent the conditional distribution of $Y\mid X$, and the black dots indicate model predictions $\hat{y}^{(k)}$. The balls indicate calibrated prediction sets $\mathcal{C}^{(k)}(x;\hat{\alpha}(z))$.
@@
We highlight the necessity of using generative models in estimating (12) by contrasting them with traditional point prediction models commonly used in standard CP. As illustrated in Figure 4.2.3, point prediction models may produce outputs that lie near the boundary of, or even outside, the inverse feasible region $\pi^{-1}(z)$, leading to overly conservative risk estimates (i.e., extremely small or even zero values when predictions fall outside the region). In contrast, generative models enable multiple stochastic draws; increasing $K$ improves the likelihood that at least one sample lies within $\pi^{-1}(z)$, yielding more accurate and less conservative risk estimates. This insight will later be formalized through the notion of the true positive rate in our theoretical analysis. The complete algorithmic procedure is summarized in Algorithm 1.
@@ 5 Theoretical Analysis
In this section, we now establish the statistical properties of our risk estimator $\hat{\alpha}(z)$, providing three main theoretical results. First, we show that our estimator is “conservative” (Theorem 5.1), ensuring it provides a valid upper bound on the true decision risk. Second, we prove that the estimator is asymptotically consistent (Proposition 5.3), admitting a tight characterization for the decision risk in large samples. Finally, we highlight the pivotal role of the generative model in our framework by analyzing the estimator’s true positive rate (Proposition 5.4), which directly influences the quality of downstream decisions. All proofs are deferred to Appendix 10.
@@ 5.1 Validity
A fundamental requirement for any risk assessment tool is that it provides reliable guarantees. We therefore first establish that CREDO’s risk estimates are conservative for both the $e$- and $p$-value radii, ensuring practitioners can trust the bounds provided by our framework.
@@
Under Section 4.1 and $R(\alpha)=R_{e}(\alpha)$, the estimator $\hat{\alpha}(z)$ in (12) satisfies
@@
Thus, $\hat{\alpha}(z)$ provides an expectation-wise upper bound on the true decision risk, without requiring correctness of the generative model. This highlights the robustness of the $e$-value construction. We now examine the corresponding guarantee for the $p$-value radius.
@@
Under Section 4.1 and $R(\alpha)=R_{p}(\alpha)$,
@@
The additional TV term reflects the deviation from full exchangeability caused by post-hoc selection of $\hat{\alpha}(z)$ (barber2023conformal, gauthier2025backward). While $R_{e}$ enjoys exact conservativeness, we later show that $R_{p}$ offers improved empirical tightness (or accuracy).
@@ 5.2 Consistency
We next study whether $\hat{\alpha}(z)$ converges to the true risk under mild assumptions on the generative model, measured through total variation distance $d_{\rm TV}(\cdot,\cdot)$.
@@
Let $\hat{\mathcal{P}}_{\mathcal{Y}|\mathcal{X}}$ be the conditional distribution learned by the generative model $g$. If for some $\delta\geq 0$,
$d_{\rm TV}\left(\hat{\mathcal{P}}_{\mathcal{Y}|\mathcal{X}=x},\mathcal{P}_{\mathcal{Y}|\mathcal{X}=x}\right)\leq\delta\quad\text{for all }x\in\mathcal{X},$
@@
Proposition 5.3 guarantees marginal consistency as $K\to\infty$, provided that the generative model approximates the conditional distribution within TV distance $\delta$. Thus, with a well-trained model and sufficiently many generated samples, $\hat{\alpha}(z)$ converges at a parametric rate to the true decision risk, complementing the finite-sample conservativeness of Theorem 5.1.
@@ 5.3 True Positive Rate
We conclude by analyzing how often the estimator correctly retains feasible decisions. Define the true positive rate as
$\mathrm{TPR}(K)\coloneqq\frac{\mathbb{E}\left[\#\{z\in\mathcal{Z}:\alpha(z)<1~\text{and}~\sum_{k=1}^{K}\tilde{\alpha}^{(k)}(z)<1\}\right]}{\#\{z\in\mathcal{Z}:\alpha(z)<1\}},$ (13)
where $\alpha(z)=\mathbb{P}\{z\notin\pi(Y)\}$ is the true risk. TPR measures the fraction of genuinely feasible decisions that are not mistakenly discarded by the estimator. From a practical perspective, TPR quantifies the ability of CREDO to correctly identify decisions that have non-trivial optimality probability while maintaining conservative guarantees. A low TPR corresponds to more “false-positive” exclusions, which may eliminate profitable or optimal decisions; thus, a higher TPR indicates better decision quality.
@@
Together, these results demonstrate that CREDO provides conservative risk estimates (validity), converges to the true risk under reasonable conditions (consistency), and benefits from generative sampling to avoid excessive conservatism (high TPR). These properties ensure both theoretical rigor and practical utility in decision support applications.
@@ 6 Computational Implementation
The theoretical analysis in the previous section establishes that CREDO provides conservative risk estimates with strong finite-sample guarantees. However, realizing these guarantees in practice requires efficiently evaluating (11) for each decision $z$ and generated sample $\hat{y}$. This evaluation presents a fundamental computational challenge as we must verify whether a calibrated conformal ball $\mathcal{C}(x;\alpha)$ is entirely contained within the inverse feasible region $\pi^{-1}(z)$. This is a set containment problem that is NP-hard in general.
This section develops practical algorithms to address this challenge. We begin by reformulating the geometric containment constraint into a tractable optimization problem (Section 6.1). Subsequently, we present efficient solution strategies for two important problem classes: linear programs, where we derive closed-form expressions, and general convex problems, where we employ gradient-based approximations (Section 6.2). These computational tools enable practitioners to deploy CREDO for real-world decision support while maintaining the theoretical guarantees established in Section 5.
6.1 Set-Containment Constraint Reformulation
@@
where $R^{-1}:[0,\infty)\to[0,1]$ is the inverse radius function:
@@
The inverse functions for the three calibrated radii in Equations (8)–(10) admit closed-form expressions, which makes the evaluation of the optimization problem explicit.
@@
Namely, the risk estimator is essentially a Monte Carlo estimator of whether the generated samples from the generative model $f(x)$ fall inside $\pi(\hat{y}^{(k)})$.
@@ 6.2.1 Linear Problems
We first examine the special case in which the decision problem is a linear program of the form
@@
where $A\in\mathbb{R}^{m\times d}$ and $b\in\mathbb{R}^{m}$ define a nonempty polyhedral feasible region. Let $\mathcal{V}$ denote the set of extreme points of this polytope, which can be solved from $A$ and $b$ using procedures such as the double description method (motzkin1953double, fukuda1995double).
In this setting, we can derive a closed-form solution for the optimization problem (14b) by leveraging $\mathcal{V}$, utilizing the property that the optimal solution to a linear program always occurs at an extreme point (bertsimas1997introduction). Specifically, by the linearity of $f(z;y)$, the constraint $f(z;y)>\min_{z^{\prime}\in\mathcal{Z}}f(z^{\prime};y)+\epsilon$ is equivalent to requiring that for all extreme points [^1] Excluding $z$ because it never satisfies the constraint with $\epsilon\geq 0$. $v\in\mathcal{V}\setminus\{z\}$, there is $\langle y,z\rangle>\langle y,v\rangle+\epsilon,$ which can be rearranged as a linear inequality $\langle y,z-v\rangle>\epsilon.$ Consequently, its feasible region is the union of a finite number of halfspaces, each associated with a direction vector zvz-v for $v\in\mathcal{V}\setminus\{z\}$. Therefore, the original optimization problem (14b) reduces to evaluating a finite set of closed-form expressions, yielding an explicit formula for the risk estimate.
Proposition 6.4 (Closed-Form Risk for Linear Programs)
Under (18), the risk estimate $\tilde{\alpha}(z)$ for a single generated sample $\hat{y}$ admits the closed-form expression:
@@ 6.2.2 Convex Problems
We adopt a heuristic algorithm (Algorithm 2) that combines ideas from the difference-of-convex algorithm (DCA) (tao1997convex) and coordinate descent (wright2015coordinate) to solve for Problem (14b). By the monotonically decreasing property of $R^{-1}(\cdot)$, we begin by noticing that the original optimization problem can be reformulated as first solving the following single-level minimization problem [^2] Without loss of generality, we assume $\mathcal{Y}$ to be the whole space.
@@
In our experiments, we empirically validate this adopted approach by showing it is reliable for producing high-quality risk estimates, providing a computationally tractable solution to decision risk assessment tasks in broader convex settings.
@@
4:  $\tilde{f}^{(t)}(z;y)\leftarrow f(z;y^{(t-1)})+\left\langle\nabla_{y}f(z;y)\big|_{y=y^{(t-1)}}~,~y-y^{(t-1)}\right\rangle.$
@@ 7 Experiments
In this section, we evaluate the performance of CREDO through extensive numerical experiments. Our results demonstrate that: ($i$) High-quality risk estimation (Section 7.1). CREDO produces risk estimates that can be tuned toward either conservativeness or accuracy through the choice of radius, with both aspects outperforming baseline models. This flexibility underscores its suitability for decision risk-assessment tasks. ($ii$) Risk-aware decision prescriptions (Section 7.2). The risk estimates produced by CREDO can be directly used to prescribe decisions that achieve consistent low risk, demonstrating CREDO’s reliability as a practical tool for risk-aware decision making. (iiiiii) Effective modular design (Section 7.3). Each component of CREDO exhibits superior performance on its respective subtask compared with alternative ablation variants, highlighting the advantages of our modeling choices.
Our experiments span a variety of optimization structures, including linear programming (LP), quadratic programming (QP), second-order conic programming (SOCP), integer programming (IP), and a penalized knapsack problem that uses semi-synthetic data from a real-world grid infrastructure investment problem (Real Data). Specifically, LP includes two stylized settings referred to as Setting I and Setting II, shown in Figure 7. Setting I features a triangular feasible region with three vertices, with a highly stochastic distribution of $Y$. Setting II represents a more complex scenario with an octagonal feasible region comprising five vertices, with a more dispersed and multimodal distribution of $Y$. These two settings are consistently used throughout the experiments due to their high interpretability and simplicity. The variance of the synthetic data distribution is controlled by a $\sigma$ parameter, which is specified by the user. Additional details of optimizations are provided in Section 11.
@@ 7.1 Risk Estimation Evaluation
Specifically, we consider five representative baseline models: ($i$) SA: sample average of the historical data $\epsilon$-optimality condition indicators; ($ii$) LR: logistic regression; (iiiiii) NN: a neural network-based (multi-layer perceptron) classifier; ($iv$) QR: quantile regression model; ($v$) CP: conformal prediction model using the upper prediction interval as its output. These baseline models capture the types of prediction approaches that decision-makers are likely to adopt as main candidates for estimating decision risk. We adopt two evaluation metrics for the decision risk estimate: ($i$) Validity: the ratio of trials where the estimated risk successfully upper bounds the true risk; ($ii$) MAE: the mean absolute error of the risk estimate compared to the true risk. These two metrics represent quantifies conservativeness and accuracy, which are the two crucial dimensions of risk estimate evaluation. Additional details of the experimental settings are included in Section 11.
Section 7.1 summarizes the results [^3] Results are reported as mean $\pm$ standard deviation across trials. Boldface indicates the best result, and underlining indicates the second best within the same setting. This convention applies to all tables in the experiment.. For the first three baselines (SA, LR, NN), we observe low estimation error but poor validity. This is expected: these models are trained to maximize predictive accuracy, not conservativeness, and thus systematically underestimate risk. In contrast, the last two baselines (QR and CP) exhibit higher validity but substantially larger errors. Although both approaches can enforce conservativeness by adjusting the quantile level $q$, the choice of $q$ is continuous in $[0,1]$ and lacks any principled mechanism to guarantee validity across problems or distributional shifts. By comparison, CREDO achieves the best of both worlds. With $p$-value and $e$-value radii, CREDO attains $100\%$ validity across all settings; with the Monte Carlo radius, it yields highly accurate risk estimates with (near) minimal error. Together, these results highlight the versatility of CREDO: it can be tuned to deliver either guaranteed conservativeness or high-precision estimation, offering a principled and robust framework for decision risk assessment.

### v3

@@ 4.2.3 Proposed Algorithm
where the radius $R(\alpha)$ is obtained by calibrating the nonconformity scores $\{L_{i}\}_{i=1}^{n}$ on $\mathcal{D}$. Of note, $R(\cdot)$ must be a decreasing function, meaning that a smaller $\alpha$ corresponds to a larger $R(\alpha)$ and higher empirical coverage. We consider three specifications of the radius function $R(\alpha):[0,1]\to\mathbb{R}_{+}\cup\{\infty\}$, each by default satisfies the boundary conditions $R(\alpha)=\infty$ for $\alpha\in[0,1/(n+1))$ and $R(1)=0$. For $\alpha\in[1/(n+1),1)$, define:
@@
$R_{\infty}(\alpha)$ $=\infty.$ (Monte Carlo radius) (9)
We note that $R_{p}$ corresponds to the classical conformal radius based on the empirical quantiles of calibration residuals (vovk2005algorithmic, singh2024distribution); $R_{e}$ is an $e$-value–based variant (grunwald2024safe), offering stronger post-hoc validity guarantees (vovk2025conformal, balinsky2024enhancing, gauthier2025values); and $R_{\infty}$ is used in a Monte Carlo-based estimation. We will discuss their respective properties in Section 5 and Section 7, showing that $R_{e}$ provides the strongest robustness, $R_{\infty}$ the highest accuracy, and $R_{p}$ a balanced trade-off between the two.
@@
Comparison of risk estimation using a point model and generative models ($K=3$ and $K=4$) in CREDO. The gray shaded area represents the inverse feasible region $\pi^{-1}(z)$. The blue dotted ellipsoids represent the conditional distribution of $Y\mid X$, and the black dots indicate model predictions $\hat{y}^{(k)}$. The balls indicate calibrated prediction sets $\mathcal{C}^{(k)}(x;\hat{\alpha}(z))$.
@@
We highlight the necessity of using generative models in estimating (11) by contrasting them with traditional point prediction models commonly used in standard CP. As illustrated in Figure 4.2.3, point prediction models may produce outputs that lie near the boundary of, or even outside, the inverse feasible region $\pi^{-1}(z)$, leading to overly conservative risk estimates (i.e., extremely small or even zero values when predictions fall outside the region). In contrast, generative models enable multiple stochastic draws; increasing $K$ improves the likelihood that at least one sample lies within $\pi^{-1}(z)$, yielding more accurate and less conservative risk estimates. This insight will be formalized through the notion of the true positive rate in our theoretical analysis (Section 5). The complete algorithmic procedure is summarized in Algorithm 1.
@@ 5 Theoretical Analysis
In this section, we now establish the statistical properties of our risk estimator $\hat{\alpha}(z)$, providing three main theoretical results. First, we show that our estimator is conservative (Theorem 5.1 and Corollary 5.2), ensuring it provides a valid upper bound on the true decision risk. Second, we prove that the estimator is asymptotically consistent (Proposition 5.3), admitting a tight characterization for the decision risk in large samples. Finally, we highlight the pivotal role of the generative model in our framework by analyzing the estimator’s true positive rate (Proposition 5.4), which directly influences the quality of downstream decisions. All proofs are deferred to Appendix 9.
@@ 5.1 Validity
A fundamental requirement for any risk assessment tool is that it provides reliable guarantees. We therefore first establish that CREDO’s risk estimates are conservative for both the $e$- and $p$-value radii, ensuring its trustworthiness to the practitioners.
@@
Under Section 4.1 and $R(\alpha)=R_{e}(\alpha)$, the estimator $\hat{\alpha}(z)$ defined in (11) satisfies
@@
Thus, $\hat{\alpha}(z)$ provides an expectation-wise upper bound on the true decision risk, without requiring correctness of the generative model. This highlights the robustness of the $e$-value construction. On the other hand, the $p$-value radius has the following guarantee.
@@
Under Section 4.1 and $R(\alpha)=R_{p}(\alpha)$, the estimator $\hat{\alpha}(z)$ defined in (11) satisfies
@@
Here, the additional TV term reflects the deviation from full exchangeability caused by post-hoc selection of $\hat{\alpha}(z)$ (barber2023conformal, gauthier2025backward), which disrupts the conservativeness of $p$-value radius’s risk estimate but was circumvented by $e$-value radius. However, we show in our experiments (Section 7.1) that conservativeness for $p$-value radius does hold in practice, and its accuracy is also superior compared to $e$-value radius. This implies that the additive term in Corollary 5.2 is usually small in practice, making $p$-value radius a more effective choice when a theoretical guarantee is not the top priority.
@@ 5.2 Consistency
We next study whether $\hat{\alpha}(z)$ converges to the true risk under mild assumptions on the generative model, measured through total variation distance $d_{\rm TV}(\cdot,\cdot)$. The result is summarized as follows.
@@
Let $\hat{\mathcal{P}}_{Y|X}$ be the conditional distribution learned by the generative model $g$. If for some $\delta\geq 0$,
$d_{\rm TV}\left(\hat{\mathcal{P}}_{Y|X=x},\mathcal{P}_{Y|X=x}\right)\leq\delta\quad\text{for all }x\in\mathcal{X},$
@@
Proposition 5.3 guarantees marginal consistency as $K\to\infty$, provided that the generative model approximates the conditional distribution within TV distance $\delta$. Thus, with a well-trained generative model and sufficiently many generated samples, $\hat{\alpha}(z)$ converges at a parametric rate to the true decision risk, complementing the finite-sample conservativeness of Theorem 5.1.
@@ 5.3 True Positive Rate
Finally, we analyze how often the estimator correctly retains feasible decisions. Define the true positive rate as
$\mathrm{TPR}(K)\coloneqq\frac{\mathbb{E}\left[\#\{z\in\mathcal{Z}:\alpha(z)<1~\text{and}~\hat{\alpha}(z)<1\}\right]}{\#\{z\in\mathcal{Z}:\alpha(z)<1\}},$ (12)
where $\alpha(z)=\mathbb{P}\{z\notin\pi(Y)\}$ is the true risk. TPR measures the fraction of genuinely feasible decisions that are not mistakenly discarded by the estimator. From a practical perspective, TPR quantifies the ability of CREDO to correctly identify decisions that have non-trivial optimality probability while maintaining conservative guarantees. A low TPR corresponds to more “false-positive” exclusions, which may eliminate profitable or optimal decisions; thus, a higher TPR indicates better decision quality. We have the following result.
@@
Together, these results demonstrate that CREDO provides conservative risk estimates (validity), converges to the true risk under reasonable conditions (consistency), and benefits from generative sampling to avoid excessive conservativeness (high TPR). These properties ensure both theoretical rigor and practical utility in decision support applications.
@@ 6 Computational Implementation
The theoretical analysis in the previous section establishes that CREDO provides conservative risk estimates with strong finite-sample guarantees. However, realizing these guarantees in practice requires efficiently solving (10) for each decision $z$ and generated sample $\hat{y}$. This presents a fundamental computational challenge as we must verify whether a calibrated conformal ball $\mathcal{C}(x;\alpha)$ is entirely contained within the inverse feasible region $\pi^{-1}(z)$, which is NP-hard in general.
This section develops practical algorithms to address this challenge. We begin by reformulating the set containment constraint into a standard optimization problem (Section 6.1). Subsequently, this allows us to derive efficient solution strategies for two important problem classes (Section 6.2): ($i$) linear programs (LPs), where we derive closed-form expressions, and ($ii$) general convex problems, where we employ gradient-based approximations. These computational insights enable practitioners to deploy CREDO for real-world decision support while maintaining the theoretical guarantees established in Section 5. The proofs in this section are deferred to Section 9.
6.1 Reformulating the Set-Containment Constraint
@@
where we expand the notation $\pi$ to $\pi_{\epsilon}$ to avoid ambiguity, and $R^{-1}(\cdot)$ is the inverse radius function:
@@
Furthermore, the inverse radius functions for the three calibrated radii in Equations (7)–(9) of Proposition 6.1 admit closed-form expressions, which makes solving for (13b) a more explicit task.
@@
Additionally, we note that these results also shed light on the namesake of the Monte-Carlo radius $R_{\infty}$, which is summarized in the following remark.
@@
Namely, the risk estimator is essentially a Monte Carlo estimator of whether the generated samples from the generative model $g(x)$ fall inside $\pi(\hat{y}^{(k)})$.
@@ 6.2.1 Linear Problems
We first examine the special case in which the decision problem is a linear program (LP) of the form
@@
where $A\in\mathbb{R}^{m\times d}$ and $b\in\mathbb{R}^{m}$ define a nonempty polyhedral feasible region. We use $\mathcal{V}$ to denote the set of extreme points of this polytope, which can be solved from $A$ and $b$ using procedures such as the double description method (motzkin1953double, fukuda1995double).
In this setting, we can derive a closed-form solution for the optimization problem (13b) by leveraging $\mathcal{V}$ and the property that the optimal solution to an LP always occurs at an extreme point (bertsimas1997introduction). Specifically, by the fact that optimal solutions for LPs are always incurred at extreme points (krein1940extreme), the constraint in (13b) can be equivalently expressed as
$f(z;y)>\min_{z^{\prime}\in\mathcal{Z}}f(z^{\prime};y)+\epsilon\quad\Longleftrightarrow\quad\text{For all }v\in\mathcal{V}\setminus\{z\},\quad\langle y,v-z\rangle-\epsilon<0,$
Consequently, its feasible region is the union of a finite number of halfspaces, each associated with a direction vector zvz-v and offset $\epsilon$. Therefore, the nested optimization problem (13b) reduces to evaluating a finite set of closed-form expressions, yielding an explicit formula for the risk estimate.
Corollary 6.4 (Closed-Form Risk Estimation for Linear Programs)
Under (17), the risk estimate $\tilde{\alpha}(z)$ for a single generated sample $\hat{y}$ admits the closed-form expression for (13b):
@@
where we also expand the notation $\pi$ to $\pi_{\epsilon}$ to avoid ambiguity.
@@ 6.2.2 Convex Problems
We adopt a heuristic algorithm (Algorithm 2) that combines ideas from the difference-of-convex algorithm (DCA) (tao1997convex) and coordinate descent (wright2015coordinate) to solve for Problem (13b). By the monotonically decreasing property of $R^{-1}(\cdot)$, we begin by noticing that the original optimization problem can be reformulated as first solving the following single-level minimization problem
@@
In Section 7.3, we empirically validate this adopted approach by showing it is reliable for producing high-quality risk estimates, providing a computationally tractable solution to decision risk assessment tasks in broader convex settings.
@@
4:  $\tilde{f}^{(t)}(z;y)\leftarrow f(z;y^{(t-1)})+\left\langle\nabla_{y}f(z;y)\big|_{y=y^{(t-1)}}~,~y-y^{(t-1)}\right\rangle$;
@@ 7 Experiments
In this section, we evaluate the performance of CREDO through extensive numerical experiments. Our results demonstrate that: ($i$) High-quality risk estimation (Section 7.1): CREDO produces risk estimates that can be tuned toward either conservativeness or accuracy through the choice of radius, with both aspects outperforming baseline models. This flexibility underscores its suitability for decision risk-assessment tasks. ($ii$) Risk-aware decision prescriptions (Section 7.2): The risk estimates produced by CREDO can be directly used to prescribe decisions that achieve consistent low risk, demonstrating CREDO’s reliability as a practical tool for risk-aware decision making. (iiiiii) Effective modular design (Section 7.3): Each component of CREDO exhibits superior performance on its respective subtask compared with alternative ablation variants, highlighting the advantages of our modeling choices. Additional details of the experiments are provided in Section 10.
We test CREDO across a diverse set of optimization problems, including linear programming (LP), quadratic programming (QP), second-order conic programming (SOCP), integer programming (IP), and a penalized knapsack problem that uses semi-synthetic data from a real-world grid infrastructure investment problem (Real Data). Specifically, we consider two stylized setups for the LPs, referred to as Setting I and Setting II, as illustrated in Figure 7. Setting I features a triangular feasible region with three vertices and a highly stochastic distribution of $Y$. Setting II introduces a more complex octagonal feasible region with five vertices and a more dispersed, multimodal distribution of $Y$. These two settings serve as running examples throughout our experiments due to their clarity and simplicity. The variance of the synthetic data distribution is controlled by a hyperparameter $\sigma$. Tabular results report the mean $\pm$ standard deviation across all trials. Boldface indicates the best-performing method, and underlining indicates the second-best within each setting.
@@ 7.1 Risk Estimation Evaluation
Specifically, we consider five representative baseline models: ($i$) SA: sample average of indicator functions computed from historic data; ($ii$) LR: logistic regression; (iiiiii) NN: a neural network-based (multi-layer perceptron) classifier; ($iv$) QE: quantile estimator; ($v$) CP (papadopoulos2002inductive, vovk2005algorithmic): standard conformal prediction model using the upper prediction interval as its output. These baseline models capture the types of prediction approaches that decision-makers are likely to adopt as main candidates for estimating decision risk. We adopt two evaluation metrics for the decision risk estimate: ($i$) Validity: the ratio of trials where the estimated risk successfully upper bounds the true risk; ($ii$) MAE: the mean absolute error of the risk estimate compared to the true risk. These two metrics quantify conservativeness and accuracy, which are the two crucial dimensions of risk estimate evaluation. Additional details of the experimental settings are included in Section 10.
Section 7.1 summarizes the results. For the first three baselines (SA, LR, NN), we observe low estimation error but poor validity. This is expected: these models are trained to maximize predictive accuracy, not conservativeness, and thus systematically underestimate risk. In contrast, the last two baselines (QE and CP) exhibit higher validity but substantially larger errors. Although both approaches can enforce conservativeness by adjusting the quantile level $q$, the choice of $q$ is continuous in $[0,1]$ and lacks any principled mechanism to guarantee validity across problems or distributional shifts. By comparison, CREDO achieves the best of both worlds. With $p$-value and $e$-value radii, CREDO attains $100\%$ validity across all settings; with the Monte Carlo radius, it yields highly accurate risk estimates with (near) minimal error. Together, these results highlight the versatility of CREDO: it can be tuned to deliver either guaranteed conservativeness or high-precision estimation, offering a principled and robust framework for decision risk assessment.